OpenStudy (anonymous):

i have to solve 2x-1/x+8 + 34x/x squared - 64

OpenStudy (anonymous):

Can you type that out in equation?

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

Is this your equation ?\[\frac{2x-1}{x+8}+\frac{34x}{x^2-64}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

There is no equal sign, so you can't exactly solve it. But think what can you multiply one of the fractions to make it look like the other so you can simply add them.

OpenStudy (anonymous):

Are you trying to find out what x is or are you trying to simplify the equation

OpenStudy (anonymous):

simplify them. and i got the common denominator as x squared - 64 but the book says that the answer is the denominator being x-8

OpenStudy (anonymous):

Maybe you can factorize \[x^2 -64\]

OpenStudy (anonymous):

once you find a common denominator you should get: \[\frac{(2x-1)(x-8)+34x}{x^2-64}\] \[x^2-64=(x+8)(x-8)\] once you expand the numerator and factor the denominator you should get: \[\frac{2x^2+17x+8}{(x+8)(x-8)}\] now you can factor the numerator: \[\frac{(2x+1)(x+8)}{(x+8)(x-8)}\] The x+8 in the numerator and denominator can be canceled giving: \[\frac{2x+1}{x-8}\]